OpenStudy (anonymous):

A particle moving along the x axis in simple harmonic motion starts from its equilibrium position, the origin, at t = 0 and moves to the right. The amplitude of its motion is 2.60 cm, and the frequency is 1.30 Hz. (f) Determine the total distance traveled between t = 0 and t = 1.15 s. PLEASE HELP!

OpenStudy (mos1635):

x=Asinωt A=2.6 cm ω=2*π*f=2*3.14*1.3=8.164 rad/sec at t=1.15 sec partical will be at position x=2.6 sin 8.164*1.15 x=2.6 sin10.79689 x=2.6 * 0.18732799 x=0.487 cm now from f and t we calculate N=number of full oscilations f=N/t N=f*t N=1.3*1.15 N=1.495  total distance= 4A+A+(A-x)=6A-x=6*2.6-0.487=15.113cm